### v3

# Evidence of Memory Effects in the Dynamics of Two-Level System Defect Ensembles Using Broadband, Cryogenic Transient Dielectric Spectroscopy

Abstract
Two-level system (TLS) defects in dielectrics cause decoherence in superconducting circuits, yet their origin, frequency distribution, and dipole moments remain poorly understood. Current probes, primarily based on qubits or resonators, require complex fabrication and measure defects only within narrow frequency bands and limited mode volumes, restricting insight into TLS behavior in isolated materials and interfaces. We introduce Broadband Cryogenic Transient Dielectric Spectroscopy (BCTDS), a broadband 3D waveguide technique that enables probing of TLS ensembles at cryogenic temperatures. Complementary to the dielectric dipper method, this approach probes a broader spectrum and reveals interference of drive-induced sidebands in TLS ensembles. The broadband, power-tunable nature of BCTDS makes it well suited for studying dressed-state physics in driven TLS ensembles, including multi-photon processes and sideband-resolved dynamics. By analyzing Fourier-transformed time-domain signals, BCTDS reveals eigen-mode frequencies of undriven TLS ensembles through characteristic V-shaped features and uncovers memory effects arising from interactions and broadband excitation. The modular method can be applied throughout device fabrication, informing mitigation strategies and advancing the design of low-loss materials with broad implications for quantum technologies and materials science.
I Introduction
TLS defects are necessary to describe anomalous low-temperature properties of disordered solids, and they play a critical role in limiting the coherence of quantum technologies such as superconducting qubits and resonators. At low temperatures, many amorphous solids exhibit a thermodynamic behavior that deviates significantly from the predictions of Debye theory, which describes the specific heat of crystalline materials as scaling with . In contrast, experiments in the early 1970s revealed that glassy materials display a nearly linear temperature dependence of the specific heat at cryogenic temperatures [1]. To describe this low temperature behavior, Anderson, Halperin, and Varma [2], and independently Phillips [3], introduced the standard tunneling model, illustrated in Fig. 1(a). The model treats disordered materials as a collection of effective double-well potentials where quantum tunneling can occur. Today, it is widely believed that individual TLS defects are ubiquitous and form an ensemble whose collective low-energy excitations account for the observed linear temperature dependence of the specific heat capacity. Early models treated TLS defects as non-interacting and uniformly distributed in frequency, which sufficed to explain the specific heat capacity measurements. However, studies have shown that TLS defects can interact, leading to spectral diffusion, clustering, and nonlinear loss [4, 5, 6, 7, 8, 9, 10].
While the notion of a TLS defect was initially developed to explain the thermodynamic properties of glassy materials, such defects have recently regained relevance in solid-state quantum computing technologies. This renewed interest is especially evident in superconducting circuits, which have become a leading platform for quantum computing primarily due to their fast gate times, design flexibility, and compatibility with advanced semiconductor fabrication techniques. However, their decoherence times are relatively short compared to other promising quantum technologies, such as neutral atoms, trapped ions, and spin qubits. From measurements of bulk loss tangents [12, 13], and modeling of the Purcell effect [14, 15, 16], it has become clear that other sources of loss must be present. Furthermore, measurements of qubit coherence while changing electric field and mechanical strain provide explicit evidence of TLS defects in superconducting circuits [17, 18, 19, 11, 20]. Currently, the prevailing understanding is that decoherence in superconducting circuits is dominated by TLS defects on surfaces and interfaces, as illustrated in Fig. 1(b), although their specific atomistic origin remains unclear [21, 22, 23, 24, 25].
TLS defects are typically probed via photonic or acoustic resonators [26, 27, 28, 29, 30] and qubits [12, 31, 17, 18, 19, 11]. These approaches provide high cooperativity but preferentially detect TLS defects with dipole moments aligned with the directly integrated probes, and are restricted to small bandwidths and mode volumes, as depicted in Fig. 1(c). Despite these limitations, several studies have demonstrated the existence of TLS-TLS interactions [32, 33, 19], and non-Markovian dynamics [34, 35], leading to questions about the nature of TLS defect interactions in the absence of strongly coupled probes [36]. In the spirit of the Mollow triplet, our setup enables the observation of dressed-state features in interacting ensembles of TLS defects, where the underlying spectral structure can reflect many-body dynamics rather than single-emitter coherence [37]. Furthermore, our method can be extended to accurately map the frequency and dipole moments of ensembles of TLS defects by employing a polarization-dependent drive. While conventional techniques such as electron spin resonance (ESR) focus on detecting net magnetization changes associated with spin-active defects using narrowband field excitation, BCTDS enables parallel, broadband interrogation of both spin and non-spin defects by directly measuring their transient electromagnetic and dielectric responses.
In addition to their impact on the decoherence of superconducting qubits, TLS defects also influence the fidelity of qubit readout. In particular, TLS defects can manifest in Josephson traveling wave parametric amplifiers (JTWPAs), which are crucial for high-fidelity qubit readout due to their quantum-limited amplification and several GHz of bandwidth [38, 39]. However, recent experiments show that high-power pulses can excite long-lived dielectric echoes in JTWPAs, which are attributed to ensembles of TLS defects [40, 41]. These echoes impair the readout fidelity of experiments utilizing JTWPAs (and likely other parametric amplifiers), motivating careful characterization of the transient dynamics of interacting ensembles of TLS defects.
Here, we propose a modular, non-invasive technique that enables BCTDS of ensembles of TLS defects in isolated materials and devices. Through systematic and longitudinal studies, this technique will provide essential data that will shed light on the atomistic nature of TLS defects and their properties [42]. Such insights will inform the design of the next generation of quantum hardware from a first-principles understanding of TLS defects, something which was recently highlighted as a pressing need for the superconducting qubit community by the Google Quantum AI hardware team [43].
II BROADBAND CRYOGENIC TRANSIENT DIELECTRIC SPECTROSCOPY
II.1 Modular, Broadband 3D Microwave Waveguide
To probe the dielectric response in the microwave frequency domain, we leverage a broadband 3D waveguide, shown in Fig. 2(a). In contrast to qubit and resonator-based probes, this method allows for coupling to ensembles of TLS defects over a broad frequency range. As described in Appendix A, the closed waveguide comprises two WR-229 to coaxial cable adapters joined by a sample clamp. This modular approach allows us to easily measure samples of different sizes, shapes, and material compositions across a broadband range of 3–6 GHz. The corresponding S-parameter curves are shown in Fig. 2(b). Importantly, these features distinguish our approach from existing narrow-band qubit and resonator-based probes, which require samples to go through a full fabrication cycle, followed by wire bonding and packaging, making it nearly impossible to perform additional material processing and characterization. Furthermore, this constraint makes the development of new fabrication recipes wasteful and inefficient.
II.2 Dielectric Response of an Ensemble of TLS Defects
Applying an external time-dependent electric field to dielectric materials induces a polarization that reflects the collective dielectric response. This response is dominated by a distribution of atomic-scale TLS defects at low temperatures, whose energy splittings and coupling strengths vary widely due to structural disorder. Within the linear response regime, the relationship between the applied field and the induced polarization is captured by the dynamical susceptibility of the material, which encodes both in- and out-of-phase components of the ensembles of TLS defects’ response. In our experiment, the response can extend beyond the linear regime when the system is influenced by strong, broadband, or pulsed electric fields, making the susceptibility dependent on amplitude, leading to higher-order contributions to the susceptibility. In this context, we use the dynamical susceptibility framework as a conceptual tool to interpret the response, even though saturation and non-linear effects likely play a significant role. After the drive is turned off, the system enters a transient regime, where the energy absorbed by TLS defects begins to dissipate. By monitoring the subsequent time evolution of the polarization-dependent homodyne signal, we perform BCTDS, which reveals signatures of the TLS defects dynamics.
To understand this system, we first consider a single TLS defect characterized by an energy detuning and a tunneling amplitude , as illustrated in Fig. 1(a). The effective Hamiltonian for the -th TLS defect in the position basis can be written as [19]
where and are Pauli matrices in the position basis representing the difference between left and right localized states.
Diagonalizing this Hamiltonian yields the energy eigenstates with eigenenergies
and defines a mixing angle through
Transforming to this energy eigenbasis, the static Hamiltonian simplifies to
where now acts in the energy eigenbasis.
Next, we introduce the interaction of the ensembles of TLS defects with an external, time-dependent electric field . The coupling occurs via the electric dipole moment of the TLS defect [37]. Choosing the dipole moment direction as the quantization axis (here aligned along ), the collective polarization operator for an ensemble of TLS defects can be defined as
where the Pauli operators are in the energy eigenbasis.
The full Hamiltonian, including the interaction with the electric field, is then written as
In our experimental setup, the electromagnetic field is polarized and can be represented as
where is the complex amplitude and the driving frequency.
This formulation captures the static energy structure of the ensembles of TLS defects and its collective coupling to the external, oscillating electric field. However, in practice, the dynamical response includes additional features such as Stark shifts and nonlinearities [44]. Unlike prior narrow-band experiments that probe TLS defects indirectly using a qubit or resonator, our approach directly drives the entire ensemble of TLS defects. This motivates including the drive term in Eq. 6 within the standard tunneling model.
The interaction in Eq. 6 causes an induced dipole moment with expectation value . Assuming that the medium is isotropic and the polarization is parallel to the electric field, we can express the average polarization using the Kubo formula [45]
where is the amplitude of the electric field, and the susceptibility is given by
where denotes the equilibrium expectation value and is the Heaviside step function. Note that the Kubo framework introduced here is only a conceptual tool to motivate the spectroscopic technique. The observed transient response in our experiment seems to suggest nonlinear contributions to the susceptibility, and we are likely far from the linear regime. However, we can still proceed by considering the Fourier transform of the induced polarization, which is given by
where
Here we will be primarily concerned with the loss induced by such an interaction, which is given by the imaginary component of Eq. 11
When the electromagnetic field interacts with the insulating sapphire samples, the loss described by Eq. 12 causes a decrease in the measured transmitted signal. However, when this drive is turned off and we enter the transient regime, the energy absorbed from the drive must now be re-emitted and ultimately measured from the output signal of the superconducting waveguide. To extract the radiated field, we perform homodyne detection, which involves extracting the classical output field quadratures and and computing the amplitude . By input-output theory [46], the output field must be proportional to the dipole response,
Thus, following Novotny et. al. [47], we can express the imaginary part of susceptibility, , using the experimentally obtained intensity as
II.3 Experimental Implementation
In principle, BCTDS can be performed on any dielectric sample at any temperature. Furthermore, even metallic or superconducting samples with oxide layers can be probed in reflection (Appendix D, Fig. 9). Here, we will be primarily interested in the dielectric response of TLS defects, which solely arise at cryogenic temperatures, as described in Sec. I. Therefore, to extract information about the TLS defects, we perform BCTDS on our samples by inserting them into the waveguide and cooling them to 10 mK using a dilution refrigerator. As mentioned in the previous section, driven dielectric spectroscopy monitors the response during the pulse, while transient dielectric spectroscopy monitors the post-pulse region (also known as the ring-down).
To probe a driven dielectric response, we generate square pulses in the 3-5 GHz band, similar to those employed for superconducting-qubit gate operations. The simple square pulses we explore here are synthesized directly on an FPGA (Field Programmable Gate Array) board (AMD RFSoC 4x2) using the Quantum Instrumentation Control Kit (QICK) software package [48]. We send the pulses through a room-temperature amplifier (Minicircuits CMA-83LN+) and into a dilution refrigerator (Bluefors LD400) where they are attenuated and ultimately reach the 3D waveguide and interact with the sample (See setup diagram in Appendix F). The resultant signal is amplified using cryogenic high-electron-mobility transistor amplifiers from Low Noise Factory before being measured on the RFSoC.
Following the cessation of the external drive, the induced polarization in the medium relaxes, leading to a coherent emission in the transient dielectric response, which imprints the temporal decay of the ensembles of TLS defects on the outgoing field, allowing us to perform BCTDS. A strong time-dependent drive may also modify the phonon-induced decay rate of a TLS defect by hybridizing its states (dressed states), thereby shifting transition frequencies and enabling decay through induced sidebands. This may alter the coupling to the spectral density of the phonon bath, leading to either enhanced or suppressed relaxation depending on the drive strength. Although we do not investigate this effect here, the broadband nature of our technique should enable its detailed study in future experiments.
III Results
III.1 Cryogenic Dielectric Spectroscopy of Different Materials
As described in Sec. II.1, our waveguide enables a modular characterization of the driven and transient dielectric response of ensembles of TLS defects in various samples. The samples considered in this study started with a 50.8 mm diameter, 432 m thick, wafer of ultra-high-purity HEMEX sapphire from Crystal Systems, which has been measured to have bulk loss tangents of the order using the dielectric dipper technique [13]. We spin coat the samples with photoresist and dice them into rectangular chips with 27.5 5.5 mm dimensions. We clean the samples with solvents following a standard procedure described in Appendix B and mount them in the waveguide.
Figure 3 summarizes the dielectric spectroscopy results obtained for different materials. We perform room-temperature control experiments on an empty waveguide and on a waveguide with bare sapphire samples that serve as benchmarks, shown in Fig. 3(a) and 3(b), respectively. Given that the TLS defects observed in our experiments have frequencies between 3 and 5 GHz, small temperature changes can significantly impact their thermal occupation, a key signature that has been explored in previous work [49, 50]. Furthermore, increases in temperature cause thermal occupation of phonon modes, which can lead to incoherent shaking of the double-well potential governing the TLS defects, likely broadening or destroying their spectral features. Our room temperature measurements in Fig. 3(aii) and 3(bii) confirm this thermal saturation effect.
As motivated in Sec. II.3, when we cool down to cryogenic temperatures and remove this thermal effect, linear response theory suggests that the spectral dips of the driven dielectric response should be corroborated by the longer-lived coherence times in the transient region. This is consistent with our findings when we cool down the same setups in Fig. 3(a) and 3(b) to 10 mK, we see an apparent narrowing of the spectral dips in the driven dielectric spectroscopy and a corresponding increase in the coherence of the transient dielectric response, shown in Fig. 3(c) and 3(d). These features are clear signatures of coherent energy storage and re-emission by the ensembles of TLS defects, which we discuss further in the context of master equation and Floquet theory in the next section (and with further investigations in Appendix G). In Fig. 3(c), we observe defect features in the absence of mounted samples, which we attribute to oxide layers present on the waveguide and antenna surfaces, and uncontrolled modes in the dilution fridge enclosure that couple to the waveguide. The signal in Fig. 3(d) likely originates from similar sources but may also include contributions from the solvent-cleaned sapphire samples in the waveguide, where residual surface contamination can result from leftover photoresist, adventitious carbon, or surface and dicing damages. In future work, we intend to remove these baseline contributions by retracting the samples in situ with a piezoelectric positioner, inspired by the dielectric dipper approach [13]. This calibration will enable further exploration of surface treatments such as etching, annealing, and passivation to reduce the measured transient dielectric response of bare sapphire and silicon wafers, which form the basis of all superconducting circuits [51, 52, 53].
Today, most Josephson junctions are fabricated using various versions of double-angle evaporation, where the insulating Josephson junction layer is formed from a native oxide grown on deposited aluminum. However, all regions of the chips that contain aluminum will inadvertently grow such native oxides, and therefore, most superconducting circuits inevitably contain large areas of aluminum oxide. To emulate the native oxide in Josephson junctions and on aluminum metal layers, we grow a thin, 2 nm aluminum oxide layer in an atomic layer deposition tool on a bare sapphire wafer. The results of these dielectric spectroscopy measurements are shown in Fig. 3(e). Here, we see clear signatures of ensembles of TLS defects, primarily in the transient region. The transient response exhibits pronounced collapse and revivals due to the interplay between external driving and TLS-defect ensemble interactions, clearly indicating memory effects [11, 54]. Furthermore, we want to highlight that the deposited 2 nm of aluminum oxide is approximately times the volume of the substrate, but is composed of the same atomic species. These findings demonstrate the importance of material morphology in the existence and behavior of TLS defects and motivate future work on passivating metallic layers [55] and developing crystalline insulating layers for Josephson junctions. This work also suggests why tantalum, which hosts a thin, stoichiometric native oxide, is an advantageous material choice for high-coherence superconducting circuits [51, 25].
Finally, we also characterize the dielectric response of 1-4 m of photoresist on sapphire, as shown in Fig. 3(f). This investigation is especially relevant because most resonators and waveguides leverage photolithography, meaning residual photoresist may be left on substrates if improperly removed. These photoresists are complex polymer chains that likely contain a high density of TLS defects [56, 57] (a detailed description of sample preparation is included in Appendix B). By understanding the dielectric properties of the resist layer, we can better assess its impact on junction performance and guide fabrication and device design optimizations [43]. Therefore, in Fig. 3(f), we observe more transient response, consistent with our hypothesis that photoresist has many TLS defects.
The technique presented in this manuscript also enables future longitudinal studies where the sample can be characterized at each step of the fabrication process, and it envisions active mitigation strategies for these defects in the next generation of superconducting qubit devices. It is essential to highlight that in the data presented in Fig. 3, we observe that the transient emission processes have lifetimes on the order of ns. Previous work on so-called dielectric echoes in JTWPA experiments [40, 41] measured lifetimes of s. Here, since the waveguide provides a large density of states and probes large ensembles, the lifetimes of these TLS defects may be reduced due to Purcell enhancement and stimulated emission. While the lifetimes extracted in these measurements are relatively short compared to state-of-the-art coherence times for superconducting qubits, the observed transient response can still play a prominent role in larger quantum circuits where gates are sequentially implemented. This effect may partially explain why large quantum circuits have overall fidelities lower than a naïve calculation based on single and two-qubit gate fidelities extracted using randomized benchmarking [58].
III.2 Dielectric Response under Different Drive Lengths
In order to better examine the relationship between absorption in the driven dielectric response and the transient emission, we zoom into a narrower frequency range of 3.9-4.4 GHz and vary the pulse lengths in Fig. 4(a)-4(c), featuring 20 ns, 50 ns, and 200 ns pulses. Increasing the drive length improves the spectral resolution and reveals finer resonant features, which manifest themselves as sharper absorption dips during the pulse. After the drive is turned off, a similar effect is seen in the transient emission, where we observe sharper and more prolonged ringdowns. Without controlled access to the physical parameters of the TLS defects, it is encouraging to observe changes in the TLS dynamics as a result of different pulses. These findings support the feasibility of pulse-based coherent control of ensembles of TLS defects.
To explore the connection between spectral dips and the surrounding emission features, we performed numerical simulations of two interacting driven spins. The full pulse evolution was treated using a master equation, while the periodic portion of the drive was analyzed separately using Floquet formalism, as detailed in Appendix G and illustrated in Fig. 12. During the pulse, the system is dressed by the external periodic drive, resulting in the emergence of sidebands. The corresponding Floquet quasi-energies show clear dips in Fig. 12(b), and prolonged ring-downs within the vicinity of these dips, as seen in Fig. 12(a). In Fig. 13, we qualitatively reproduce the sharpening of ring-downs for different pulse durations as observed in the experiment.
III.3 Further Investigation of Cryogenic Transient Dielectric Spectroscopy Results
In this section, we provide a more in-depth analysis of the transient dielectric response from the sample in Fig. 3(f) containing photoresist on sapphire, which is shown in Fig. 5. We start by plotting the transient dielectric response in Fig. 5(aii) showing long-timescale oscillatory collapse and revival behavior of the emitted signal, which indicates information backflow and memory effects [59, 60]. This can be seen clearly in Fig. 5(ai), which displays the response at a single frequency of GHz. To quantify the frequency components of the oscillations observed in the time traces in Fig. 5(a), we perform a FFT of all the data traces to generate Fig. 5(bii). Again, a particular line cut at the same frequency is shown in Fig. 5(bi). The FFT shows clear V-shaped features and horizontal bands, particularly in the 3.65–4.0 GHz and 4.25–4.65 GHz ranges. These V-shapes arise from off-resonant emission by dominant eigenfrequencies of the ensembles of TLS defects, made visible by the broadband response of the waveguide [61]. Numerical simulations of a driven, interacting four-spin system (Appendix G, Fig. 13) confirm this behavior, namely the appearance of V-shaped arcs in the FFT, which sharpen with longer pulse durations, and converge toward the bare eigenmode frequencies.
We also compute the two-time correlation function [62]
where is the intensity of the transmitted pulse, and plot it in Fig. 5(c). provides us with a map of temporal correlations and displays evident non-exponential, oscillatory decay, which is a hallmark signature of memory effects or non-Makovian dynamics of the ensembles of TLS defects. Finally, we utilize the computed to extract using Eq. 14, which we show in Fig. 5(d). Here, consistent with , we observe non-monotonic and negative , highlighted by the cyan contours at zero in Fig. 5(d), exhibiting partial overlap with the aforementioned V-shapes in some regions of the FFT. This behavior indicates physics beyond linear response and further supports the presence of memory effects and re-emission in the ensembles of TLS defects [63, 64].
IV Conclusions and Outlook
This work introduces BCTDS, a technique to probe TLS defects that supports fast and modular mounting of various samples, without invoking narrow-band probes like resonators or qubits. Our measured cryogenic dielectric spectroscopy experiments reveal distinct transient features and memory effects as demonstrated by the and plots in Fig. 5(c), 5(d), respectively. Such memory effects may indicate that the dynamics are non-Markovian, which may be detrimental for quantum error correction if manifested in quantum processors [65, 66, 54]. We believe that much of the memory effects arises from the interplay between coherent TLS defect interactions and driving and measuring in a broadband environment. By analyzing the frequency and temporal correlations of large ensembles of TLS defects in isolated materials, we take initial steps toward uncovering the atomistic structure of TLS defects. Next, we aim to introduce an orthogonal SMA port to rotate the polarization of the drive and perform the same dielectric spectroscopy experiments, subject to different drive strengths. By integrating over different drive polarizations, we can better understand the frequency distribution and dipole moments of the TLS defects in various samples.
While this work has focused on materials for superconducting qubits, we highlight that BCTDS is a more general technique that provides information about the loss of various dielectric materials. This study is also not exclusively applicable to TLS defects; for instance, we can use BCTDS to measure vacancies and boron dopants in silicon and nitrogen vacancies in diamond directly, without the need for optical interrogation [67]. In conclusion, we have demonstrated a novel spectroscopy tool based on a broadband 3D waveguide that enables global interrogation of ensembles of TLS defects. The approach is entirely modular, and we demonstrate that this can be leveraged for longitudinal devices that can help develop causal relationships between properties of ensembles of TLS defects and material processing steps.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Acknowledgements.
Bert Harrop graciously diced the devices at Princeton. Sofia Abrego and Chloe Buschmann helped review and polish the manuscript. We also acknowledge stimulating conversations with Kater Murch, Miles Blencowe, Geoffroy Hautier, Chandrasekhar Ramanathan, and Valla Fatemi. Startup funds from the Thayer School of Engineering at Dartmouth primarily supported this work. While the scientific focus of this work is quite distinct from existing funding, M.F. would like to express gratitude for the support of science within his group through the DARPA Young Faculty Award No. D23AP00192 and NSF PHY-2412555, and hopes that projects like this continue to be made possible by grants from the US federal government. The views and conclusions in this manuscript are those of the authors. They should not be interpreted as representing the official policies, expressed or implied, of DARPA, the NSF, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.Author Contributions
M.F. proposed the waveguide approach and guided the experimental and theoretical work. Q.W. performed the experiments, made the figures, and developed the data analysis pipeline for this work. S.M.G. and J.F. designed and fabricated the waveguide. J.S.S.G., R.L., and S.B. provided theoretical support. J.S.S.G., S.B., and M.F. provided context for the work. S.B. provided theoretical insights using an interacting-driven TLS model to validate the observed patterns. All authors jointly validated the results and participated in the writing of the manuscript.
Competing Interests
The authors declare no competing interests.
Additional information
Correspondence and request for materials
Correspondence and material requests should be directed to Mattias Fitzpatrick (mattias.w.fitzpatrick@dartmouth.edu). Data supporting the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Waveguide Design and Fabrication
A.1 Waveguide Design
The fundamental building block of our waveguides is an adapter that provides a 50 impedance-matched conversion between an SMA coaxial cable and a WR-229 3D rectangular waveguide. To understand the basic features of the waveguide, we solve the wave equation for a rectangular prism, with boundary conditions that the electric field goes to zero at the walls of the waveguide. After performing separation of variables, we can identify a set of allowed eigenmodes for the cross section perpendicular to the direction of propagation, with wavenumbers given by
where and are the width and height of the rectangular cross-section of the waveguide, respectively. The constants and are integers whose combinations identify different eigenmodes of the cross-section. This means that the propagation constant of the transmitted wave can be expressed as
where is the wavenumber. Significantly, Eq. 17 predicts that when , becomes imaginary, and the solutions become decaying exponentials along the waveguide in a domain known as the waveguide’s cutoff [14]. Therefore, we can describe the frequencies of the allowed modes of the waveguide using
where is the speed of light. From Eq. 18, we then identify the cutoff frequency, which is the lowest frequency mode, given by
assuming that . The cutoff frequency formula provided in Eq. 19 is idealized and assumes a perfect rectangular waveguide, which is challenging to manufacture in practice due to the sharp inner corners. Therefore, we utilize Ansys HFSS to consider the design of the actual waveguide in this work. The drive and readout ports are coaxial SMA cables that transition into small antennas integrated into the waveguide. This coupling scheme smoothly transforms the coaxial mode of the SMA cables to a close approximation the mode of an ideal rectangular waveguide, producing a polarized electric field that will ultimately couple to the TLS defects in the sample.
Figure 7 provides detailed dimensions of our SMA to WR-229 3D rectangular waveguide adapter. In this work, we leverage Eq. 19 to achieve the desired 3 to 6 GHz frequency band and adapt the perfect rectangular waveguide designs to make the manufacturing easier. In particular, our design features an inner radius of the 3D waveguide designed to use a 3/8 inch end mill. M2.5 screws are used to mount the 4.1 mm SMA pins that integrate into the designs shown in Fig. 7.
We fabricate two copies of the SMA to WR-229 3D rectangular waveguide adapter to construct a full waveguide. Then, we construct a clamp for samples that can be introduced between the two waveguide adapters. This approach allows us to consider a variety of sample layouts and sizes. The assembly of the complete waveguide can be seen in Fig. 6.
A.2 Waveguide Fabrication
The waveguide used in this manuscript is manufactured from Aluminum 6061 using a HAAS Super Mini Mill in Dartmouth’s Thayer School of Engineering machine shop. A combination of steel wire wool and polishing stones was used to improve the surface finish, especially the rounded corners, and remove any feed marks left behind from the end mill. Finally, the waveguide was cleaned using isopropanol (IPA) to remove impurities and oily residues.
Appendix B Sample Processing
To demonstrate the modularity of our platform, we consider three types of samples, the preparation of which is explained in this section.
We start with high-grade crystalline 0.5 mm-thick sapphire wafers from Crystal Systems, which are then diced into 27.5 x 5.5 mm rectangular strips, the standard substrate size for hosting 3D transmons. The samples are initially covered in resist AZ 1518 to protect them from debris during dicing (likely another source of TLS defects). We soak the chips in acetone at 70 ∘C overnight and sonicate them in acetone before transferring them to IPA and drying with compressed nitrogen gas.
To study the effects of photoresist on the TLS density, we spin-coat Shipley 1813 with thicknesses ranging from 1 to 4 m. We also study the oxide layers (typically a few nanometers in thickness), another candidate source of TLS defects. Instead of evaporated aluminum, which would give both an aluminum and oxide layer, we used an atomic layer deposition (ALD) machine to grow a 2 nm thin film. We expect the entire film to be oxidized by doing 20 burst cycles at 80 ∘C and repeatedly using H2O as an oxidant at each burst.
Appendix C Cryogenic Dielectric Spectroscopy of the Same Sample, For Different Cooldowns
Two-level system (TLS) defects are known to shift in frequency and spatial location after thermal cycling [68], due to rearrangements in the disordered environment that reshape the local energy landscape. To corroborate this signature, we perform dielectric spectroscopy experiments 15 days apart on the same sample subject to a thermal cycle to 300 K in Fig. 8. As expected, we see a significantly different driven and transient dielectric response in Fig. 8(a) and 8(b), highlighted by the difference plot in Fig. 8(c). In future work, we intend to perform systematic studies of the temperature dependence closer to 10 mK. Such studies may provide insight into the observed ‘universal dielectric response’, and potentially probe the transition of quantum TLS defects to the broad classical dielectric response [69, 70].
Appendix D Dielectric Response in Transmission and Reflection
For ideal, lossless systems, transmission and reflection contain equivalent information as incident power must be conserved. In our experiment, where dissipation is present, reflection provides complementary information about dielectric loss during the pulsing region. We make measurements on the same samples as in Appendix C but place a cryogenic circulator at the waveguide input to measure both the transmission (Fig. 9(a)) and reflection (Fig. 9(b)) signals. In the transient region, emission from the dielectric samples is detected almost equally in the transmission and reflection channels, as the samples are positioned equidistant from the two antenna pins, and the readout circuits are identical. As a result, the collected signal shows minimal difference, as shown in Fig. 9(c). A short vertical white line is added to indicate pulse bandwidth f.
Appendix E Electric Field Amplitude Calibration
The waveguide design ensures that the transmitted signal is relatively constant over a bandwidth of approximately 3-6 GHz, shown in Fig. 2(b). However, the samples under test are coupled in the middle of the waveguide, which can be frequency-dependent, as illustrated in Fig. 10(a). We leverage HFSS simulations to predict the average electric field amplitude at the sample plane over our frequency bandwidth , plotted in blue dots and lines in Fig. 10(b). To compensate for this variation, we apply a frequency-dependent scaling factor (red solid line). The impact of this calibration is evident when comparing the transient spectra of a sapphire sample with an AlOx deposit before (Fig. 10(c)) and after (Fig. 10(d)) applying the correction. Significantly, most of the key features in the data remain. Still, such calibrations can ensure uniform electric field densities as a function of frequency and may prove critical in future experiments to deduce properties of individual ensembles of TLS defects.
Appendix F Measurement Setup
This work utilizes an AMD radio-frequency system on a chip (RFSoC) board, which provides a fast sampling rate (9.85 GSPS for DAC and 5 GSPS for ADC) and direct synthesis and readout of microwave pulses without needing up- and down-conversion. To control the RFSoC, we utilize the Quantum Instrumentation Control Kit, a Python package developed by Fermi National Accelerator Laboratory [71].
The pulses generated from the RFSoC pass through a room-temperature amplifier (Minicircuits CMA-83LN+) and a series of attenuators at various temperature stages of a Bluefors LD400 dilution refrigerator. The signal reaches the waveguide input port at base temperature (typically 10 mK) and interacts with the sample (Fig. 11). The output signal from the waveguide is collected using a readout SMA pin that then passes through a cryogenic circulator, HEMT (High Electron Mobility Transistor) amplifier, and two room-temperature amplifiers (Minicircuits CMA-83LN+), before returning to the RFSoC. A detailed wiring diagram is provided in Fig. 11.
Appendix G Qualitative Comparison to Numerical Simulations
The purpose of the numerics is to qualitatively reproduce the key experimental observation that the Fourier spectra exhibit broad sidebands for short pulses, which progressively sharpen as shown in Fig. 4 in the main text. To this end, we model the system using a driven-spin model with a finite-duration pulse, which provides a simplified framework for interpreting the interference patterns observed in the TLS ring-down and understanding how the applied drive shapes the dynamics. During this interval, the system is expected to exhibit Floquet-like physics, with the dynamics governed by the quasienergy spectrum and associated sidebands. After the end of the pulse, signatures of the prior Floquet dynamics may persist in the post-pulse evolution, manifesting as coherent oscillations or ringdown features. To support this interpretation, we briefly summarize the Floquet formalism and illustrate the emergence of sidebands using numerical simulations. We first briefly summarize the formalism following [72, 73]. Consider a time-periodic Hamiltonian with period ,
and define the Floquet frequency .
Floquet theorem states that solutions to the time-dependent Schrödinger equation are as follows.
can be expressed as
where the Floquet mode is periodic,
and is the quasi-energy of Floquet.
Expand into Fourier harmonics,
leading to the eigenvalue equation in the extended space
where
For a monochromatic drive
the Floquet Hamiltonian is block-tridiagonal:
The dipole operator defines the polarization. The dipole response function is given by the retarded commutator.
where is the Heaviside step function and is the expectation in the steady-state Floquet. Using Floquet states, the response can be expanded as
where the Fourier components satisfy the Dyson equation
Explicitly, the Fourier transform of the dipole response function is expressed via the Floquet Green’s functions, as
where are the elements of the Floquet dipole matrix, , and the poles
determine the resonance frequencies.
When the system is driven by a periodic square pulse of duration , then turned off at , the system freely evolves with no drive for . The post-pulse time evolution contains oscillatory terms at frequencies set by quasi-energy differences and sideband shifts , weighted by the residues of these poles in . These produce ring-down oscillations, whose decay rates reflect the imaginary parts of .
For a driven two-level system with detuning and external field amplitude , the time-dependent Hamiltonian is given by
Here, the periodic drive induces dressed states with quasi-energies
which hybridize near multi-photon resonances .
The effective coupling strength of photons is proportional to where is a Bessel function of the first kind. This effective coupling changes the size of the avoided crossing near multi-photon resonance. These avoided crossings form characteristic V-shaped structures near resonances in the plane, where the arms of the V correspond to bare energy levels shifted by multiples of , and the vertex marks the opening of a gap due to multi-photon coupling. The tree-like structures observed in the ringdowns arise from the interference of multiple sideband emissions. Unlike the case of a single driven two-level system, our system consists of an ensemble of interacting TLSs coupled via dipole-dipole interactions. As a result, each eigenstate of the bare, time-independent Hamiltonian becomes dressed by the external drive, leading to more complex interference effects that manifest as beatings in the observed signal.
Although the discussion above focuses on a single two-level system for illustrative purposes, our numerical simulations consider systems with and coupled spins, incorporating weak dipole-dipole interactions, which are three orders of magnitude smaller than the individual TLS frequencies. Using this model, we observe collapse and revival patterns that qualitatively resemble the experimental observations. The Fourier transform of the simulated dynamics reveals characteristic V-shaped features in the spectral domain, converging at the bare eigenfrequencies of the coupled system. Dissipation is modeled through a collective radiative decay channel using the Lindblad master equation, which may be justified in our setup due to the presence of a broadband waveguide mediating the system–environment interaction. Hence, including interactions and collective dissipation, the Hamiltonian becomes [19]
The dynamics is governed by the Lindblad master equation.
where are collective lowering and raising operators. The assumption of collective decay is motivated by the broadband nature of the waveguide, which couples similarly to all TLSs.
Figure 12 shows numerical simulations of the collective excitation dynamics and Floquet quasi-energies in a two-spin system. Panel (a) displays the time evolution of the collective population, obtained by solving the Lindblad master equation with fixed dipole-dipole coupling, drive amplitude, pulse duration (marked by a horizontal dashed line), collective dissipation, and two distinct TLS frequencies (indicated by vertical dashed lines). During the drive, clear Rabi-Chevron patterns emerge, while following the pulse, pronounced interference and long-lived ring-down oscillations are observed. Figure 12(b) shows the Fourier amplitude of the full time-domain simulation as a function of drive frequency. The V-shaped features converging toward the bare TLS frequencies reflect detunings between the drive and the individual transitions. These structures highlight the coherent excitation conditions that give rise to the long-lived ring-downs observed in panel (a). Figure 12(c) displays the first four Floquet quasi-energies computed during the driving period. The quasi-energy spectrum exhibits characteristic dips near the driving frequencies, which correspond to resonances that enhance coherent excitation trapping. These resonances manifest as long-lived ring-down dynamics in the time domain, indicating that the system temporarily stores energy in quasi-stationary Floquet states with suppressed decay, thus prolonging the coherent oscillations observed after the pulse ends.
Figure 13 presents numerical simulations of a driven, interacting four-spin network under a finite-duration pulse. In the experiment, the emitted signal is measured via homodyne detection of the in-phase (I) and quadrature (Q) components, which are proportional to the collective dipole radiation from the system. This output field corresponds to the collective lowering operator , so the measured power reflects the expectation value , directly connecting to the simulated observable . The upper panels show the time-domain evolution of this collective excitation as a function of drive frequency for increasing pulse durations, where pronounced interference patterns emerge in the ring-down following the pulse, as shown in Fig. 13(a). Figure 13(b) displays the corresponding Fourier spectra, which reveal how the spectral content of the ring-down evolves with increasing pulse duration. For short pulses, the abrupt temporal cutoff introduces large energy uncertainty, resulting in broad sidebands. As the pulse duration increases, the drive becomes more spectrally selective, allowing coherent oscillations among near-resonant modes to build up over time. This temporal coherence leads to narrower sidebands in the Fourier spectrum, which progressively sharpen and approach the bare transition frequencies of the system. This sharpening gives rise to the observed V-shaped structure in the frequency domain that we observe experimentally.
The theoretical framework used here shows good qualitative agreement with the experimental data, capturing key features such as prolonged ring-downs that sharpen with increasing pulse duration, leading to more defined spectral features. This is particularly notable given the general assumptions underlying the model, especially in our TLS framework, where driving is incorporated within the standard tunneling model alongside dipole–dipole interactions. These encouraging results motivate future work aimed at a more detailed theoretical and numerical investigation of the atomistic origin and universal dielectric response of the ensembles of TLS defects probed by our broadband transient spectroscopy technique.